OpenStudy (anonymous):

Maria is 3 years older than her brother, Luis. Luis is 2 years older than their younger sister, Karla. The sum of their ages is 55. How old are the three siblings?

OpenStudy (anonymous):

L+M+K = 55 K+2=L L+3=M Do you know how to solve systems?

OpenStudy (anonymous):

m.. If I do I dont remember? I've done this problem before, I just lost my paper so I have to do it all over again. I think I did something like x+5=55 because I was starting with Karlas age and then when i got that I just added 3 and two to those numbers to get Luis and Maria's ages. would that work?

OpenStudy (anonymous):

Its a good attempt but you need more than one variable since youve got 3 unknowns.

OpenStudy (anonymous):

Aw dang. ): Then no I have no idea how to solve systems